# Conformal e-prediction

### v3

@@ Abstract
This paper discusses a counterpart of conformal prediction for e-values, conformal e-prediction. Conformal e-prediction is conceptually simpler and had been developed in the 1990s as precursor of conformal prediction. When conformal prediction emerged as result of replacing e-values by p-values, it seemed to have important advantages over conformal e-prediction without obvious disadvantages. This paper re-examines relations between conformal prediction and conformal e-prediction systematically from a modern perspective. Conformal e-prediction has advantages of its own, such as the ease of designing conditional conformal e-predictors and the guaranteed validity of cross-conformal e-predictors (whereas for cross-conformal predictors validity is only an empirical fact and can be broken with excessive randomization). Even where conformal prediction has clear advantages, conformal e-prediction can often emulate those advantages, more or less successfully. Conformal e-prediction can also serve as basis for testing; the resulting conformal e-testing looks very different from but inherits some strengths of conformal testing.
@@ 1 Introduction
In fact, e-values were used (under different names) when discussing precursors of conformal prediction in the 1990s; in this paper we will refer to this variety of conformal prediction as conformal e-prediction. One early description of conformal e-prediction is [5], and [1] is a recent rediscovery. The paper [27] that first introduced conformal prediction also discusses conformal e-prediction. In this paper, we will occasionally refer to conformal prediction as conformal p-prediction in order to emphasize it being based on p-values.
Soon after the publication of [5, 27], conformal e-prediction seems to have disappeared. Perhaps the main reason why it was superseded by conformal prediction was that conformal predictions can be packaged as prediction sets [28, Sect. 2.2], and in this case their property of validity is very easy to state: we just say that the probability of error is at most $\epsilon$ at a prespecified significance level $\epsilon$ [28, Proposition 2.3]. This was clearly stated only in 2001 [12, Theorem 1], although this statement was implicit in the standard requirement of validity for p-values stated in [27]. The standard requirement of validity for e-values, also stated in [27], does not admit such an intuitive restatement in terms of probability of error.
Another reason for conformal e-prediction losing its popularity may have been the finding in 2002 [23, Theorem 1] that, in the on-line mode of prediction, smoothed conformal predictors make errors independently. An important corollary of this stronger property of validity is that small probabilities of errors manifest themselves, with high probability, as a low frequency of errors [28, Corollary 2.5].
A third advantage of conformal prediction is that the strong property of validity found in [23, Theorem 1] allows us to test the assumption of exchangeability. This strong property can be stated as the independence of the conformal p-values output at different steps, and to test exchangeability, we can bet against the conformal p-values being independent and uniformly distributed. This led to the introduction in 2003 [29] of conformal test martingales. For a long time, nothing was known about the efficiency of conformal test martingales, and first results about their efficiency appeared in 2019 (see [26]); this is a major topic of [28, Part III].
The last advantage of conformal prediction that we discuss in this section was found only in 2017 [30], and so it did not contribute to the eclipse of conformal e-prediction in the early 2000s. It was the discovery of conformal predictive distributions, motivated by [19]: in the case of regression, smoothed conformal prediction produces “conformal predictive distributions”, which are automatically well-calibrated.
@@
For simplicity, in this paper we consider IID (or at least exchangeable) data and concentrate on the problem of pattern recognition (also known as classification). We start in Sect. 2 from the definition of conformal e-prediction and continue in Sect. 3 with discussing its validity. The simplest property of validity (Proposition 1 in Sect. 3) consists in conformal e-predictors producing at each step a valid e-value for the true label: namely, its expectation is at most 1 at each step; the general notion of e-value is introduced right after Proposition 1. Since the expectation is the average over the sample space, we can say that conformal e-predictors are valid in the space domain, or space-wise. A complementary notion of efficiency is efficiency in the time domain, which we consider next.
@@
Proposition 1 does not say anything about the relation between the e-values for the true labels produced at different steps. Is it possible that for some streams of data the average of the e-values produced at different steps tends to 2 while for others it tends to 0.5 with non-zero probability? In Sect. 3 we show that this is impossible in the online prediction protocol, stating both the strong law of large numbers and the law of the iterated logarithm for the e-values produced for the true labels at different steps. This does not fully replace the strong property of independence of errors for conformal prediction, but it can be regarded as a partial replacement. We can see that the e-values produced at different steps are not misleading in the time domain; not only is their expectation at most 1 at each step, their average is at most 1 time-wise in the long run.
@@
The properties of validity established in Sect. 3 are marginal, in the sense that the expectations, or time averages, in them are not conditional on any properties of the observations. For example, if our predictions are for people, in principle we can get very different averages for men and women. In the case of conformal prediction, a simple way of achieving conditional validity is using Mondrian conformal prediction [28, Sect. 4.6]. Mondrian conformal prediction requires a hard partition of observations, such as the partition of people into men and women. Interestingly, conformal e-prediction is much more flexible when trying to achieve conditional validity. Its conditional version does not have to be based on a partition (Sect. 4); e.g., we may require separate validity for men, women, and Europeans. It has been shown recently that this type of conditionality can also be achieved for conformal prediction (see, e.g., [6]), but it is much easier to achieve and appears to be more natural in the case of conformal e-prediction. This can be regarded as an advantage of conformal e-prediction.
@@
What we discuss in Sections 24 is “full” conformal e-prediction, and it is computationally inefficient for many standard ways of producing predictions. Section 5 introduces split conformal e-prediction, which is a simple way to make conformal e-prediction computationally efficient. Similarly to split-conformal prediction (introduced in [15, 14]), split conformal e-prediction can lose in predictive efficiency as compared with full conformal e-prediction.
@@
To prevent loss in computational efficiency without sacrificing predictive efficiency, cross-conformal predictors were introduced in [25]. Cross-conformal predictors are not provably valid [25, Appendix], and this sometimes even shows in experimental results [11]. The limits of violations of validity are given by Rüschendorf’s result (see, e.g., [31]): when merging p-values coming from different folds by taking arithmetic mean (this is essentially what cross-conformal predictors do), the resulting arithmetic mean has to be multiplied by 2 in order to guarantee validity. In the more recent method of jackknife+, introduced in [2] and closely related to cross-conformal prediction, there is a similar factor of 2 [2, Theorem 1], which cannot be removed in general [2, Theorem 2].
@@
In Sect. 6, we introduce a version of cross-conformal prediction based on e-values, which we call cross-conformal e-prediction. The situation with cross-conformal e-prediction is very different from cross-conformal prediction, as the arithmetic mean of e-values is always an e-value. This is an obvious fact, but it is shown in [32, Sect. 3] that arithmetic mean is the only useful merging rule. Therefore, cross-conformal e-prediction is always valid. This is a second advantage of conformal e-prediction.
@@
The emphasis of Sections 36 is on the validity of conformal e-prediction and its computational efficiency, while Sect. 7 moves on to its predictive efficiency. What are suitable criteria of predictive efficiency? We propose two such criteria in the case of pattern recognition, the “observed log criterion” and the “prior log criterion”.
In Sections 8 and 9 we will discuss various ways of testing the exchangeability assumption, one-off and “multistage”, based on conformal e-prediction. We will see that conformal e-testing inherits some strengths of conformal testing.
@@ 2 Conformal e-predictors
Suppose we are given a training set $z_{1},\dots,z_{n}$ consisting of labelled objects $z_{i}=(x_{i},y_{i})$ and our goal is to predict the label of a new object $x$. In this paper we consider predictors of the following type: for each potential label $y$ for $x$ we would like to have a number $f(z_{1},\dots,z_{n},x,y)$ reflecting the plausibility of $y$ being the true label of $x$. An example is conformal transducers [28, Sect. 2.7], which, in the terminology of this paper, may be called conformal p-predictors. The output
@@
where SV is the set of indices of support vectors: $i\in\textrm{SV}$ if and only if $z_{i}$, $i=1,\dots,n+1$, is a support vector for the SVM constructed from $z_{1},\dots,z_{n+1}$ as training set. It is based on the indicator function of being a support vector. When given a training set $z_{1},\dots,z_{n}$ and a new object $x$, this conformal e-predictor goes through all potential labels $y$ for $x$ and for each constructs an SVM and outputs $f(z_{1},\dots,z_{n},x,y)$. It makes it computationally inefficient.
@@ 3 Validity of conformal e-predictors
The following obvious proposition asserts the validity of conformal e-predictors. Let us write $Z_{1},Z_{2},\dots$ for the random vectors whose realizations are the observed data $z_{1},z_{2},\dots$; more generally, $(X,Y)$ or $Z$ are random vectors with values in the observation space $\mathbf{Z}$. As usual, a finite sequence of random vectors is exchangeable if its joint distribution does not change if it is permuted (and an infinite sequence is exchangeable if its joint distribution does not change if its first $n$ elements are permuted, for any $n$). The difference between exchangeability and being IID (independent and identically distributed) can be substantial for finite data sequences (but for infinite data sequences the difference between the two assumptions disappears according to de Finetti’s theorem [28, Sect. A.5.1]).
@@
Proposition 1 involves a “space average”, i.e., an average over the sample space. The analogous property of validity for conformal prediction can be stated in terms of error probabilities for set predictions (see, e.g., [28, Proposition 2.1], which is particularly intuitive. This is the first advantage of conformal prediction pointed out in Sect. 1. It disappears when we move to conformal e-prediction: validity has to be defined in terms of e-values (however, it has been argued [18] that e-values are more intuitive than p-values).
Another advantage of conformal prediction is that, in the online mode of prediction (to be defined shortly), conformal predictors (in their smoothed version) make errors at different steps independently (see, e.g., [28, Theorem 11.1]). Without independence, it is possible, even when the probability of error at each step is $\epsilon$, for the long-term relative frequency of errors over consecutive steps to be either 0 or 1 (1 with probability $\epsilon$). The independence of errors forces the long-term frequency of errors to be $\epsilon$ almost surely (which is also asserted in [28, Proposition 2.1]). We can say that independence implies ergodicity, i.e., the almost sure coincidence of time and space averages.
@@
To give an example demonstrating the loss of independence in conformal e-prediction, we first define the online prediction protocol. In the online protocol for conformal e-prediction, we observe an object $x_{1}$, apply the conformal e/predictor to compute the conformal e-values $f(x_{1},y)$ for all possible labels $y\in\mathbf{Y}$, observe the true label $y_{1}$, record the e-value $e_{1}:=f(x_{1},y_{1})$ for it, observe another object $x_{2}$, apply the conformal e-predictor to compute the e-values $f(x_{1},y_{1},x_{2},y)$ for all possible labels $y\in\mathbf{Y}$, observe the true label $y_{2}$, record the e-value $e_{2}:=f(x_{1},y_{1},x_{2},y_{2})$ for it, etc. In the case conformal prediction, we can instead record conformal p-values $p_{n}$ (or record whether a mistake is made at a given significance level). While the conformal p-values are independent (for smoothed conformal predictors), the following example shows that they are not already in a toy binary situation (perhaps this is the simplest non-trivial example of conformal e-prediction). We write $E_{n}$ for the e-value $e_{n}$ regarded as a random variable (an e-variable).
@@
We follow the proof of [24, Lemma 14] (given as Lemma 3.15 in the first edition of [28]) and use the terminology of [21, Chap. 7].
Let $\mathcal{F}_{n}$ be the $\sigma$-algebra generated by the multiset $\lbag Z_{1},\dots,Z_{n-1}\rbag$ and the observations $Z_{n},Z_{n+1},\dots$. (These $\sigma$-algebras form the exchangeable filtration [17, Sect. 5.6].) For each time horizon $N\in\{2,3,\dots\}$, the stochastic sequence $(E_{n}-1,\mathcal{F}_{n})$, $n=N,\dots,1$, is a bounded supermartingale difference. By Hoeffding’s inequality (see, e.g., [28, Sect. A.6.3]), for any $N\in\{2,3,\dots\}$,
@@
where $C$ is an upper bound on the conformal e-predictor $f$. By the Borel–Cantelli lemma [20, Sect. 2.10] (part (a)), the internal inequality in (7) holds only for finitely many $N$ for a fixed $\epsilon>0$. This implies (6).
@@
(where $\delta_{a}$ is the probability measure concentrated on $\{a\}$), since $E_{N},\dots,E_{1}$ are generated independently. This gives us a product distribution for $E_{1},E_{2},\dots$, namely (10) with $N$ replaced by $\infty$. For any $b\in\{1,2,\dots\}$, by Le Cam’s version of Poisson’s theorem [20, Sect. 3.12], the probability of the event that $b$ of the $E_{n}$ for $n\in[2^{k},2^{k+1})$ will satisfy $E_{n}=n$ tends to $e^{-\lambda}\lambda^{b}/b!$ as $k\to\infty$, where $\lambda:=\ln 2$; therefore, this event will happen infinitely often. Therefore, (6) will be violated almost surely; namely, the limsup in it will be $\infty$ almost surely. Figure 1 illustrates this plotting the average conformal e-value $\frac{1}{N}\sum_{n=1}^{N}e_{n}$ vs $N$ for a wide range of $N$, $N\in\{1,\dots,10^{10}\}$ (with $N$ given on a logarithmic scale).

### v4

@@ Abstract
This paper discusses a counterpart of conformal prediction for e-values, conformal e-prediction. Conformal e-prediction is conceptually simpler and had been developed in the 1990s as precursor of conformal prediction. When conformal prediction emerged as result of replacing e-values by p-values, it seemed to have important advantages over conformal e-prediction without obvious disadvantages. This paper re-examines relations between conformal prediction and conformal e-prediction systematically from a modern perspective. Conformal e-prediction has advantages of its own, such as the ease of designing conditional conformal e-predictors and the guaranteed validity of cross-conformal e-predictors (whereas for cross-conformal predictors validity is only an empirical fact and can be broken with excessive randomization). Even where conformal prediction has clear advantages, conformal e-prediction can often emulate those advantages, more or less successfully.
@@
In fact, e-values were used (under different names) when discussing a precursor of conformal prediction in the 1990s; in this paper we will refer to this precursor as conformal e-prediction. One early description of conformal e-prediction is [5]. The paper [24] that first introduced conformal prediction also discusses conformal e-prediction. In this paper, we will occasionally refer to conformal prediction as conformal p-prediction in order to emphasize it being based on p-values.
Soon after the publication of [5, 24], conformal e-prediction seems to have disappeared. Perhaps the main reason why it was superseded by conformal prediction was that conformal predictions can be packaged as prediction sets [25, Sect. 2.2], and in this case their property of validity is very easy to state: we just say that the probability of error is at most $\epsilon$ at a prespecified significance level $\epsilon$ [25, Proposition 2.3]. This was clearly stated only in 2001 [12, Theorem 1], although this statement was implicit in the standard requirement of validity for p-values stated in [24]. The standard requirement of validity for e-values, also stated in [24], does not admit such an intuitive restatement in terms of probability of error (see, however, Appendix B).
Another reason for conformal e-prediction losing its popularity may have been the finding in 2002 [21, Theorem 1] that, in the on-line mode of prediction, smoothed conformal predictors make errors independently. An important corollary of this stronger property of validity is that small probabilities of errors manifest themselves, with high probability, as a low frequency of errors [25, Corollary 2.5].
The last advantage of conformal prediction that we discuss in this section was found only in 2017 [26], and so it did not contribute to the eclipse of conformal e-prediction in the early 2000s. It was the discovery of conformal predictive distributions, motivated by [17]: in the case of regression, smoothed conformal prediction may produce “conformal predictive distributions”, which are automatically well-calibrated.
@@
For simplicity, in this paper we consider IID (or at least exchangeable) data and concentrate on the problem of pattern recognition (also known as classification). We start in Sect. 1 from the definition of conformal e-prediction and continue in Sect. 2 with discussing its validity. The simplest property of validity (Proposition 1 in Sect. 2) consists in conformal e-predictors producing at each step a valid e-value for the true label: namely, its expectation is at most 1 at each step; the general notion of e-value is introduced right after Proposition 1. Since the expectation is the average over the sample space, we can say that conformal e-predictors are valid in the space domain, or space-wise. A complementary notion of efficiency is efficiency in the time domain, which we consider next.
@@
Proposition 1 does not say anything about the relation between the e-values for the true labels produced at different steps. Is it possible that for some streams of data the average of the e-values produced at different steps tends to 2 while for others it tends to 0.5 with non-zero probability? In Sect. 2 we show that this is impossible in the online prediction protocol, stating both the strong law of large numbers and the law of the iterated logarithm for the e-values produced for the true labels at different steps. This does not fully replace the strong property of independence of errors for conformal prediction, but it can be regarded as a partial replacement. We can see that the e-values produced at different steps are not misleading in the time domain; not only is their expectation at most 1 at each step, their average is at most 1 time-wise in the long run.
@@
The properties of validity established in Sect. 2 are marginal, in the sense that the expectations, or time averages, in them are not conditional on any properties of the observations. For example, if our predictions are for people, in principle we can get very different averages for men and women. In the case of conformal prediction, a simple way of achieving conditional validity is using Mondrian conformal prediction [25, Sect. 4.6]. Mondrian conformal prediction requires a hard partition of observations, such as the partition of people into men and women. Interestingly, conformal e-prediction is much more flexible when trying to achieve conditional validity. Its conditional version does not have to be based on a partition (Sect. 3); e.g., we may require separate validity for men, women, and Europeans. It has been shown recently that this type of conditionality can also be achieved for conformal prediction (see, e.g., [6]), but it is much easier to achieve and appears to be more natural in the case of conformal e-prediction. This can be regarded as an advantage of conformal e-prediction.
@@
What we discuss in Sections 13 is “full” conformal e-prediction, and it is computationally inefficient when built on top of many standard prediction algorithms (such as neural networks). Section 4 introduces split conformal e-prediction, which is a simple way to make conformal e-prediction computationally efficient. Similarly to split-conformal prediction (introduced in [14, 13]), split conformal e-prediction can lose in predictive efficiency as compared with full conformal e-prediction.
@@
To prevent loss in computational efficiency without sacrificing predictive efficiency, cross-conformal predictors were introduced in [23]. Cross-conformal predictors are not provably valid [23, Appendix], and this sometimes even shows in experimental results [11]. The limits of violations of validity are given by Rüschendorf’s result (see, e.g., [27]): when merging p-values coming from different folds by taking arithmetic mean (this is essentially what cross-conformal predictors do), the resulting arithmetic mean has to be multiplied by 2 in order to guarantee validity. In the more recent method of jackknife+, introduced in [2] and closely related to cross-conformal prediction, there is a similar factor of 2 [2, Theorem 1], which cannot be removed in general [2, Theorem 2].
@@
In Sect. 5, we introduce a version of cross-conformal prediction based on e-values, which we call cross-conformal e-prediction. The situation with cross-conformal e/prediction is very different from cross-conformal prediction, as the arithmetic mean of e-values is always an e-value. This is an obvious fact, and it is shown in [28, Sect. 3] that arithmetic mean is the only useful merging rule. Therefore, cross-conformal e/prediction is always valid. This is a second advantage of conformal e-prediction.
@@
The emphasis of Sections 25 is on the validity of conformal e-prediction and its computational efficiency, while Sect. 6 moves on to its predictive efficiency. What are suitable criteria of predictive efficiency? We propose two such criteria in the case of pattern recognition, the “observed log criterion” and the “prior log criterion”.
@@ 1 Conformal e-predictors
Suppose we are given a training set $z_{1},\dots,z_{n}$ consisting of labelled objects $z_{i}=(x_{i},y_{i})$ and our goal is to predict the label of a new object $x$. In this paper we consider predictors of the following type: for each potential label $y$ for $x$ we would like to have a number $f(z_{1},\dots,z_{n},x,y)$ reflecting the plausibility of $y$ being the true label of $x$. An example is conformal transducers [25, Sect. 2.7], which, in the terminology of this paper, may be called conformal p-predictors. The output
@@
where SV is the set of indices of support vectors: $i\in\textrm{SV}$ if and only if $z_{i}$, $i\in\{1,\dots,n+1\}$, is a support vector for the SVM constructed from $z_{1},\dots,z_{n+1}$ as training set. It is based on the indicator function of being a support vector. When given a training set $z_{1},\dots,z_{n}$ and a new object $x$, this conformal e-predictor goes through all potential labels $y$ for $x$ and for each constructs an SVM and outputs $f(z_{1},\dots,z_{n},x,y)$. It makes it computationally inefficient.
@@ 2 Validity of conformal e-predictors
The following obvious proposition asserts the validity of conformal e-predictors. Let us write $Z_{1},Z_{2},\dots$ for the random vectors whose realizations are the observed data $z_{1},z_{2},\dots$; more generally, $(X,Y)$ or $Z$ are random vectors with values in the observation space $\mathbf{Z}$. As usual, a finite sequence of random vectors is exchangeable if its joint distribution does not change if it is permuted (and an infinite sequence is exchangeable if its joint distribution does not change if its first $n$ elements are permuted, for any $n$). The difference between exchangeability and being IID (independent and identically distributed) can be substantial for finite data sequences (but for infinite data sequences the difference between the two assumptions disappears according to de Finetti’s theorem [25, Sect. A.5.1]).
@@
Proposition 1 involves a “space average”, i.e., an average over the sample space. The analogous property of validity for conformal prediction can be stated in terms of error probabilities for set predictions (see, e.g., [25, Proposition 2.1]), which is particularly intuitive. This is the first advantage of conformal prediction pointed out in Sect. Conformal e-prediction. It disappears when we move to conformal e-prediction: validity has to be defined in terms of e-values (however, it has been argued [16] that e-values are more intuitive than p-values).
Another advantage of conformal prediction is that, in the online mode of prediction (to be defined shortly), conformal predictors (in their smoothed version) make errors at different steps independently (see, e.g., [25, Theorem 11.1]). Without independence, it is possible, even when the probability of error at each step is $\epsilon$, for the long-term relative frequency of errors over consecutive steps to be either 0 or 1 (1 with probability $\epsilon$). The independence of errors forces the long-term frequency of errors to be $\epsilon$ almost surely (which is also asserted in [25, Proposition 2.1]). We can say that independence implies ergodicity, i.e., the almost sure coincidence of time and space averages.
@@
To give an example demonstrating the loss of independence in conformal e-prediction, we first define the online prediction protocol. In the online protocol for conformal e-prediction, we observe an object $x_{1}$, apply the conformal e/predictor to compute the conformal e-values $f(x_{1},y)$ for all possible labels $y\in\mathbf{Y}$, observe the true label $y_{1}$, record the e-value $e_{1}:=f(x_{1},y_{1})$ for it, observe another object $x_{2}$, apply the conformal e-predictor to compute the e-values $f(x_{1},y_{1},x_{2},y)$ for all possible labels $y\in\mathbf{Y}$, observe the true label $y_{2}$, record the e-value $e_{2}:=f(x_{1},y_{1},x_{2},y_{2})$ for it, etc. In the case conformal prediction, we can instead record conformal p-values $p_{n}$ (or record whether a mistake is made at a given significance level). While the conformal p-values are independent (for smoothed conformal predictors), the following example shows that the conformal e-values are not independent already in a toy binary situation (perhaps this is the simplest non-trivial example of conformal e-prediction). We write $E_{n}$ for the e-value $e_{n}$ regarded as a random variable (an e-variable).
@@
We follow the proof of [22, Lemma 14] (given as Lemma 3.15 in the first edition of [25]) and use the terminology of [19, Chap. 7].
Let $\mathcal{F}_{n}$ be the $\sigma$-algebra generated by the multiset $\lbag Z_{1},\dots,Z_{n-1}\rbag$ and the observations $Z_{n},Z_{n+1},\dots$. (These $\sigma$-algebras form the exchangeable filtration [15, Sect. 5.6].) For each time horizon $N\in\{2,3,\dots\}$, the stochastic sequence $(E_{n}-1,\mathcal{F}_{n})$, $n=N,\dots,1$, is a bounded supermartingale difference. By Hoeffding’s inequality (see, e.g., [25, Sect. A.6.3]), for any $N\in\{2,3,\dots\}$,
@@
where $C$ is an upper bound on the conformal e-predictor $f$. By the Borel–Cantelli lemma [18, Sect. 2.10, part (a) of the lemma], the internal inequality in (7) holds only for finitely many $N$ for a fixed $\epsilon>0$. This implies (6).
@@
(where $\delta_{a}$ is the probability measure concentrated on $\{a\}$), since $E_{N},\dots,E_{1}$ are generated independently. This gives us a product distribution for $E_{1},E_{2},\dots$, namely (10) with $N$ replaced by $\infty$. For any $b\in\{1,2,\dots\}$, by Le Cam’s version of Poisson’s theorem [18, Sect. 3.12], the probability of the event that $b$ of the $E_{n}$ for $n\in[2^{k},2^{k+1})$ will satisfy $E_{n}=n$ tends to $e^{-\lambda}\lambda^{b}/b!$ as $k\to\infty$, where $\lambda:=\ln 2$; therefore, this event will happen infinitely often (for any $b$). Therefore, (6) will be violated almost surely; namely, the limsup in it will be $\infty$ almost surely. Figure 1 illustrates this by plotting the average conformal e-value $\frac{1}{N}\sum_{n=1}^{N}e_{n}$ vs $N$ for a wide range of $N$, $N\in\{1,\dots,10^{10}\}$ (with $N$ given on a logarithmic scale).